OpenStudy (anonymous):

solve the system xy^3=10^14 x^2/y=10^7

OpenStudy (anonymous):

y =10^3 x=10^5

OpenStudy (anonymous):

get x in terms of y from the first equation and substitute in the second equation.

OpenStudy (anonymous):

so thats the answer?